OpenStudy (anonymous):

fine the value of rel number y such that ( 3+2i)(1+iy) is (a)real and (b) imaginary

OpenStudy (michele_laino):

please you have to apply the distributive property of multiplication over addition, in order to perform your calculus

OpenStudy (michele_laino):

namely: \[(3+2i)(1+iy)=3(1+iy)+2i(1+iy)=\] please try

OpenStudy (michele_laino):

Please, I can not say you the solution directly, since the code of conduct

OpenStudy (michele_laino):

so, please continue the above calculus that I have wrote

OpenStudy (anonymous):

(3-2y)+i(2y+2) now next?

OpenStudy (michele_laino):

sorry I got: (3-2y)+i(2+3y)

OpenStudy (anonymous):

oh yes just writng mistake but now next ?

OpenStudy (michele_laino):

ok! if you want a real number, you have to set: 2+6y=0, then solve for y; similarly if you want an imaginary number, you have to set: 3-2y=0, and then solve for y again

OpenStudy (michele_laino):

sorry I have made an error, ...you have to set 2+3y=0...

OpenStudy (anonymous):

thank you but why we put os equal to zero?

OpenStudy (michele_laino):

when you set 3-2y=0, your number becomes: i(2+3y), which is an imaginary number. SImilarly, when you set 2+3y=0, your number becomes: 3-2y, which is a real number!

OpenStudy (michele_laino):

for example in the first case you get y=3/2, and your number is: i(2+3*3/2)=i*13/2 whereas in the second case you get y=... please try!

OpenStudy (michele_laino):

@ziawasim